### v1

# Worse than Random? An Embarrassingly Simple Probing Evaluation of Large Multimodal Models in Medical VQA

Abstract
Large Multimodal Models (LMMs) have shown remarkable progress in the field of medical Visual Question Answering (Med-VQA), achieving high accuracy on existing benchmarks. However, their reliability under robust evaluation is questionable. This study reveals that state-of-the-art models, when subjected to simple probing evaluation, perform worse than random guessing on medical diagnosis questions. To address this critical evaluation problem, we introduce the Probing Evaluation for Medical Diagnosis (ProbMed) dataset to rigorously assess LMM performance in medical imaging through probing evaluation and procedural diagnosis. Particularly, probing evaluation features pairing original questions with negation questions with hallucinated attributes, while procedural diagnosis requires reasoning across various diagnostic dimensions for each image, including modality recognition, organ identification, clinical findings, abnormalities, and positional grounding. Our evaluation reveals that top-performing models like GPT-4V and Gemini Pro perform worse than random guessing on specialized diagnostic questions, indicating significant limitations in handling fine-grained medical inquiries. Besides, models like LLaVA-Med struggle even with more general questions, and results from CheXagent demonstrate the transferability of expertise across different modalities of the same organ, showing that specialized domain knowledge is still crucial for improving performance. This study underscores the urgent need for more robust evaluation to ensure the reliability of LMMs in critical fields like medical diagnosis, and current LMMs are still far from applicable to those fields. https://github.com/eric-ai-lab/ProbMed
Figure 1: Accuracy of four LMMs on two types of specialized questions in medical diagnoses, with and without adversarial pairs. The significant drop in accuracy with adversarial pairs highlights the models’ unreliability in handling medical diagnoses.
1 Introduction
Foundation models, such as large language models (LLMs) (OpenAI, 2023a; Touvron et al., 2023; Jiang et al., 2023; Anil et al., 2023; Chung et al., 2022) and large multimodal models (LMMs) (OpenAI, 2023b; Team et al., 2023; Li et al., 2023b; Liu et al., 2023a; Chen et al., 2023a), have demonstrated impressive capabilities in understanding complex visual and text inputs, generating human-like language, and achieving high accuracy on various benchmarks. The integration of these foundation models into real-life medical practice holds immense potential given their advanced computational capabilities (Wu et al., 2023; Yang et al., 2023) and promising progress on existing medical Visual Question Answering (Med-VQA) benchmarks (Lau et al., 2018; Liu et al., 2021; He et al., 2020; Zhang et al., 2023). As we stand on the precipice of integrating these models into critical decision-making domains, one natural question appears: how much can we trust these models in real-world scenarios, such as medicine and healthcare, where the stakes are high?
Before discussing the reliability of LMMs in critical domains like Med-VQA, we must first address a fundamental question: Are we evaluating LMMs correctly? To address this question, we introduce a simple yet effective probing evaluation method that exposes the weaknesses of LMMs by creating simple binary questions with hallucination pairs over existing benchmarks. An example is shown in Figure 2. Despite the high accuracy reported on current Med-VQA tasks, our study reveals a significant vulnerability in LMMs when faced with adversarial questioning, as illustrated in Figure 1. The performance drops observed are alarming: even advanced models like GPT-4V and Gemini Pro perform worse than random guessing, with an average decrease of 42% across the tested models.
Figure 2: An example illustrating the potential for misleading accuracy in existing evaluations. While the model correctly identifies the position of an existing finding in the standard evaluation, it fails to differentiate between actual and hallucinated positions when subjected to an adversarial evaluation.
Based on this, we further analyze a critical question: How reliable are LMMs in medical diagnosis, ranging from general questions to specialized diagnostic questions? To address this question, we introduce ProbMed, which features procedural diagnosis designed to rigorously evaluate model performance across multiple diagnostic dimensions. We curated ProbMed from 6,303 images sourced from two widely-used biomedical datasets, MedICaT (Subramanian et al., 2020) and ChestX-ray14 (Wang et al., 2017). These images cover various modalities, including X-ray, MRI, and CT scans, and span multiple organs such as the abdomen, brain, chest, and spine. Using GPT-4 and a positional reasoning module, we generated metadata for each image, extracting information about abnormalities, condition names, and their corresponding locations. This metadata facilitated the automatic generation of 57,132 high-quality question-answer pairs, covering dimensions like modality recognition, organ identification, abnormalities, clinical findings, and positional reasoning.
Our systematic evaluation of seven state-of-the-art LMMs on ProbMed revealed several critical insights. First, even the best-performing models, such as GPT-4V and Gemini Pro, performed close to random guessing on specialized diagnostic categories like Condition/Finding and Position, highlighting their limitations in handling fine-grained medical inquiries. Second, introducing adversarial pairs significantly reduced the accuracy of all models, with LLaVA-Med’s performance dropping by up to 57% and Gemini Pro’s accuracy decreasing by 25% in ProbMed. These findings emphasize the importance of adversarial testing in Med-VQA to uncover model weaknesses. Third, the CheXagent model, which was exclusively trained on chest X-rays, demonstrated that specialized domain knowledge is crucial. It showed that expertise gained on one particular organ could be transferable to another modality of the same organ in a zero-shot manner, highlighting the value of domain-specific training for improving model performance.
In summary, our work highlights significant gaps in the reliability of LMMs for medical diagnosis despite their impressive performance on current existing general domain benchmarks. The insights from ProbMed underscore the urgent need for robust evaluation methodologies to ensure the accuracy and reliability of LMMs in real-world medical applications. This research inspires the development of more trustworthy AI systems in healthcare and beyond, ultimately contributing to better diagnostic outcomes and patient care.
2 Related Work
2.1 Large Multimodal Models in the Medical Domain
The advancements in Large Multimodal Models (LMMs) have significantly enhanced the understanding and generation of medical content that integrates both visual and linguistic elements. Notable models include GPT-4V (OpenAI, 2023b), Gemini Pro (Team et al., 2023), LLaVA (Liu et al., 2023a, b), and MiniGPT-v2 (Chen et al., 2023a). The scalability and exceptional performance of these large foundation models have driven their application in the biomedical field.
Further progress has been made in fine-tuning general-domain LMMs for the biomedical field, resulting in specialized models like LLaVA-Med (Li et al., 2023a), MedBLIP (Chen et al., 2023b), and Med-Flamingo (Moor et al., 2023). Despite the promising results from these domain-specific LMMs, there is ongoing exploration into training smaller multimodal models to address specific clinical needs. For instance, models like LLaVA-RAD (Chaves et al., 2024) and CheXagent (Chen et al., 2024) have been developed for chest X-ray interpretation, aiming to bridge competency gaps in radiology tasks. As we move towards integrating these models into critical decision-making processes, it becomes imperative to assess their reliability in high-stakes environments like healthcare and medicine.
2.2 Medical Visual Question Answering
Medical Visual Question Answering (Med-VQA) plays a crucial role in assessing the capabilities of models in interpreting and responding to queries about medical images. Some benchmarks, like VQA-RAD (Lau et al., 2018) and SLAKE (Liu et al., 2021), are manually constructed with categorical question types. While this method ensures high-quality question-answer pairs, it is labor-intensive and results in limited dataset scales.
Automated curation methods have been developed to address scalability. PathVQA (He et al., 2020) uses CoreNLP [^1] https://stanfordnlp.github.io/CoreNLP tools, and PMC-VQA (Zhang et al., 2023) employs generative models to create larger datasets. However, these methods often sacrifice fine-grained question categories, and some require additionally trained models for question filtering. ProbMed, as shown in Table 1, stands out by providing large-scale benchmarks and enabling categorical accuracy assessments across various diagnostic dimensions for each image, including modality recognition, organ recognition, clinical findings identification, and positional grounding. ProbMed uniquely incorporates adversarial negation pairs for each question-answer pair to ensure diagnostic specificity and reliability, setting it apart from existing benchmarks.
Different evaluation methods are employed for assessing LMMs, including closed-ended VQA, multiple choice VQA, and open-ended generation tasks such as captioning and report generation. Open-ended VQA and report generation are typically considered more challenging and harder to evaluate, often requiring human or model evaluation alongside automated lexical similarity metrics like ROUGE-L and BLEU-4. Recent works (Wang et al., 2024; Zheng et al., 2024) argue that multiple-choice questions may not be ideal due to inherent "selection bias." In our work, we choose a relatively easy-to-evaluate method: closed-ended VQA augmented with adversarial evaluation methods featuring hallucinated attributes. By requiring the model to accurately distinguish relevant features, we enhance the reliability of the evaluation process. This method allows for clear and definitive assessments, improving the overall robustness of our findings in medical contexts.
Table 1: Comparison ProbMed with a test set of existing medical VQA datasets, demonstrating our dataset’s difference from existing benchmarks. For SLAKE, only the English subset is considered for head-to-head comparison with existing benchmarks.
Dataset Images Questions
Question
Category
Procedural
Diagnosis
Adversarial
Pairs
VQA-RAD (Lau et al., 2018) 0.2k 0.4k
SLAKE (Liu et al., 2021) 0.09k 1k
PathVQA (He et al., 2020) 0.8k 6.7k
PMC-VQA (Zhang et al., 2023) 50k 400k
ProbMed (Ours) 6.3k 57k
3 ProbMed: Probing Evaluation for Medical Diagnosis
In this section, we design two evaluation principles and present a comprehensive analysis of state-of-the-art LMMs for Med-VQA using the created ProbMed dataset to address two research questions:
  1. 1.
    Is the current evaluation of LMMs for Med-VQA reliable?
  2. 2.
    How reliable are LMMs on medical diagnosis, ranging from general questions to specialized diagnostic questions?
Our primary goal is to rigorously evaluate these models’ readiness for real-life diagnostic tasks, particularly under adversarial conditions. Despite their high accuracy on existing benchmarks, the models struggle with simple probing evaluation. ProbMed is designed to expose these vulnerabilities and provide a more reliable assessment of model performance in real-world scenarios.
Figure 3: Flow diagram of the ProbMed data curation process. Two comprehensive biomedical datasets were utilized to collect source data and construct a metadata file, enabling the automatic generation of high-quality question-answer pairs for the ProbMed dataset.
3.1 Probing Evaluation with Adversarial Pairs
One of the main motivations behind ProbMed is to assess the models’ ability to accurately distinguish between relevant and irrelevant features. ProbMed pairs original questions with negation questions containing hallucinated attributes. This method challenges the model robustness by requiring them to identify true conditions while disregarding false, hallucinated ones. For instance, a question about a specific finding is paired with a negated question featuring a different, non-existent finding to test if the model can exclusively identify the true finding.
3.2 Procedural Diagnosis
To ensure a comprehensive evaluation, ProbMed includes questions that require reasoning across multiple diagnostic dimensions for each image. These dimensions include modality recognition, organ identification, clinical findings, abnormalities, and positional reasoning. This multifaceted approach assesses a model’s diagnostic capabilities beyond simple question-answer pairs, requiring it to integrate various pieces of information to form a coherent diagnostic picture.
3.3 Data Curation
As illustrated in Figure 3, ProbMed draws from two comprehensive biomedical datasets MedICaT and ChestX-ray14 to compile a diverse set of 6,303 images. MedICaT (Subramanian et al., 2020) contains 217k image-caption pairs from 131k open access biomedical papers. From this dataset, we selected 4,543 image-caption pairs focusing on a single organ and modality, with clear indications of normal or abnormal conditions. These images span three modalities (X-ray, MRI, and CT scan) and four organs (abdomen, brain, chest, and spine). ChestX-ray14 (Wang et al., 2017) comprises 112k frontal-view X-ray images from 30k unique patients, including 880 images with abnormalities marked by bounding boxes. We selected 1,760 images, balanced between healthy and abnormal cases, with disease labels and bounding box annotations.
To create high-quality, balanced question-answer pairs, we generated metadata for each image. For MedICaT images, GPT-4 (OpenAI, 2023a) was used to analyze captions, identify abnormalities, and extract positional descriptions using few-shot prompting. For ChestX-ray14 images, a positional reasoning module generated textual descriptors of abnormalities based on bounding boxes and image sizes. The standardized metadata across both subsets included unique condition names and positional descriptions for each organ-modality combination as candidates for later constructing adversarial question-answers.
Using the metadata, we generated a diverse set of high-quality questions for each image, covering various diagnostic dimensions. Ground-truth questions were created based on available metadata, with "yes" answers. For each ground-truth question, we also created a corresponding adversarial question by selecting random adversarial entities and assigning "no" answers. This process included selecting alternative organs, modalities, and hallucinated conditions/findings. As shown in Table 2, this process resulted in a total of 57,132 question-answer pairs, averaging 9 pairs per image.
Table 2: Dataset Statistics of ProbMed. There are 6.3k images and 57k VQA pairs in total. The dataset is balanced within each question type and image type.
Organ, Modality Image Question
Question with
Answer "yes"
Unique Condition
Unique Positional
Description
Abdomen MRI 84 757 375 107 75
Brain MRI 566 5,046 2,509 697 446
Chest MRI 40 382 189 52 38
Spine MRI 324 3,346 1,664 461 336
Abdomen CT scan 751 6,855 3,410 909 552
Brain CT scan 270 2,417 1,200 335 209
Chest CT scan 548 5,161 2,572 727 353
Spine CT scan 87 941 470 149 93
Abdomen X-ray 232 2,046 1,018 277 160
Brain X-ray 79 599 298 84 44
Chest X-ray 3,178 27,530 13,278 1,418 694
Spine X-ray 202 2,052 1,020 300 172
Total 6,303 57,132 28,003 / /
4 Experimental Analysis
We conducted a systematic evaluation and comprehensive analysis using the ProbMed dataset on seven state-of-the-art LMMs to identify their strengths and weaknesses in real-life imaging diagnostics. Apart from proprietary GPT-4V (OpenAI, 2023b) and Gemini Pro (Team et al., 2023), we selected five open-source models (7B parameters version), spanning across general models including LLaVA-v1 (Liu et al., 2023b), LLaVA-v1.6 (Liu et al., 2023a), MiniGPT-v2 (Chen et al., 2023a) and specialized models including LLaVA-Med (Li et al., 2023a) and CheXagent (Chen et al., 2024). These models were chosen based on their computational cost, efficiency, and inference speed, making them practical for integration into medical practice. For categorical accuracy, we only considered a hit if the model correctly answered all questions within the category for an image (see Table 19). This means the model must correctly identify all existing entities and disregard all hallucinated entities.
4.1 Is Current Evaluation of LMMs for Med-VQA Reliable?
To address this first research question, we introduced adversarial pairs to the evaluation process to test the model’s robustness and reliability. This strategy ensures that models must validate the absence of certain characteristics or findings rather than simply acknowledge existing conditions, thereby enhancing diagnostic specificity and reliability. To demonstrate the necessity of adversarial pairs for achieving valid and trustworthy accuracy scores in Med-VQA, we conduct an experimental analysis on the test set of an existing medical dataset, VQA-RAD (Lau et al., 2018), in addition to ProbMed.
4.1.1 Probing Evaluation with Adversarial Pairs in VQA-RAD
To construct challenging adversarial questions for an given image, ideally, we need full control over the ground truth information and a set of confusing candidates, as provided in ProbMed. However, since VQA-RAD (Lau et al., 2018) provides finalized question-answer pairs without metadata, we could only construct adversarial pairs for 118 test instances where the answer is "yes" out of 272 closed-ended question-answers pairs within its test set. Each adversarial pair was manually created such that, based on the limited information from the original question-answer pair, the answer to the adversarial question had to be negated. This process resulted in 236 question-answer pairs in total. The adversarial questions in this subset are less challenging than those in ProbMed, as they often involve a simple semantic negation of the original question due to limited information.
These results, as shown in Table 3, reveal the significant impact of adversarial pairs on model performance. Although the original accuracy appears very high for some underperforming models, the accuracy drops drastically after balancing the subset with adversarial pairs: 6.78% for GPT-4V and 16.95% for Gemini Pro, with an average decrease of 42.7% across the tested models.
Table 3: Model accuracy on the VQA-RAD test subset before and after introducing adversarial pairs. The table demonstrates the significant drop in accuracy across various models when adversarial pairs are added, highlighting the models’ vulnerabilities to adversarial questions. The percentage decrease in accuracy is noted in parentheses.
Models
Accuracy (%) without
Adversarial Pairs
Accuracy (%) with
Adversarial Pairs
LLaVA-v1-7B 96.61 25.42 (-71.19)
LLaVA-v1.6-7B 77.11 8.47 (-68.64)
LLaVA-Med 81.35 3.38 (-77.97)
MiniGPT-v2 72.03 46.61 (-25.42)
CheXagent 53.38 21.18 (-32.20)
GPT-4V 40.67 33.89    (-6.78)
Gemini Pro 61.86 44.91 (-16.95)
4.1.2 Probing Evaluation with Adversarial Pairs in ProbMed
Table 4 demonstrates the similar significant impact of adversarial pairs in ProbMed. The accuracy of more capable models is generally less affected by the introduction of challenging adversarial pairs. However, even the most robust models experience a minimum drop of 10.52% in accuracy when tested with ProbMed’s challenging questions, with an average decrease of 44.7% across the tested models, highlighting the critical role of probing evaluation in evaluating Med-VQA performance comprehensively.
Table 4: The impact of adversarial pairs on model accuracy in ProbMed. This table illustrates the performance change of models with the introduction of adversarial pairs. The results indicate a substantial decline in accuracy, underlining the robustness of ProbMed in challenging LMMs. The percentage drop in accuracy is noted in parentheses.
Models
Accuracy (%) without
Adversarial Pairs
Accuracy (%) with
Adversarial Pairs
LLaVA-v1-7B 95.40 16.51 (-78.89)
LLaVA-v1.6-7B 83.79 21.97 (-61.82)
LLaVA-Med 71.91 14.77 (-57.14)
MiniGPT-v2 85.30 22.99 (-62.31)
CheXagent 39.55 22.18 (-17.37)
GPT-4V 66.27 55.75 (-10.52)
Gemini Pro 78.36 53.26 (-25.10)
4.2 How Reliable Are LMMs in Medical Diagnosis?
After correcting model accuracy by introducing adversarial pairs, we continue to address the second research question. We conducted diagnostic probing ranging from general to specialized diagnostic questions using the ProbMed dataset.
4.2.1 Performance across Diagnostic Questions
Table 5 shows the categorical accuracy of different models aggregated among all image types. While GPT-4V and Gemini Pro outperform other models and excel in general tasks such as recognizing image modality and organs, their low performance in specialized tasks like determining the existence of abnormalities and answering fine-grained questions about condition/finding and position highlights a significant gap in their ability to aid in real-life diagnosis.
On more specialized diagnostic questions, even top-performing models like GPT-4V and Gemini Pro performed close to random guessing. Their accuracy in identifying conditions and positions was alarmingly low, underscoring their limitations in handling fine-grained medical inquiries. CheXagent, trained exclusively on chest X-rays, achieved the highest accuracy in determining abnormalities and conditions. However, its performance in general tasks like identifying image modality and organs was lower, highlighting the need for domain-specific training. LLaVA-Med achieves much higher accuracy compared to both LLaVA-v1-7B and LLaVA-v1.6-7B in identifying conditions/finding and their positions but fell short on more general tasks like identifying image modality and organs as well as determining abnormality.
Among the open-sourced general-purpose models, MiniGPT-v2 performs the best, surpassing domain-specific LLaVA-Med in identifying organs and both LLaVA-Med and CheXagent in determining positions of condition/finding without domain-specific training. A more detailed breakdown of the performance of different models on different image types across each question type is available in Appendix A.
Table 5: Categorical and overall accuracy (%) of different models aggregated among all image types in ProbMed. The best result in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice 25.00 25.00 50.00 35.67 36.48
LLaVA-v1-7B 25.28 40.53 50.00 0.34 0.10
LLaVA-v1.6-7B 6.77 80.70 46.18 3.57 1.07
LLaVA-Med 5.49 32.98 38.76 20.39 5.37
MiniGPT-v2 3.25 76.29 50.09 15.23 8.05
CheXagent 37.26 33.96 73.32 28.53 7.24
GPT-4V 92.50 71.71 53.30 35.19 22.32
Gemini Pro 96.47 75.69 62.59 27.92 17.17
4.2.2 Error Analysis in Procedural Diagnosis
For models whose accuracy dropped drastically after introducing adversarial pairs, we observed a consistent accuracy pattern much lower than random guess performance for specialized questions. An error analysis focusing on GPT-4V and Gemini Pro across three specialized question types - Abnormality, Condition/Finding, and Position is further conducted. Each accuracy measurement is conditional on the model successfully answering the preceding diagnostic questions, reflecting a procedural diagnosis approach. This analysis reveals both models’ vulnerabilities to hallucination errors, particularly as they progress through the diagnostic procedure, with Gemini Pro being more prone to accepting false conditions and positions.
As shown in Table 6, for Abnormality questions, conditioned on correctly identifying both image modality and organ, GPT-4V’s errors arise from both incorrect answers and its tendency to reject challenging questions, while Gemini Pro attained a slightly higher accuracy of 67.05%, with all errors resulting from incorrect answers.
More specialized questions in identifying conditions and their positions, conditioned on successful abnormality detection, reveal both models’ vulnerabilities to hallucination errors, particularly as they progress through the diagnostic procedure, with Gemini Pro more prone to accepting false conditions and positions. For questions on condition/finding, GPT-4V’s accuracy dropped to 36.9%, with roughly even error distribution between denying ground-truth conditions and accepting hallucinated conditions, while most of the errors of Gemini Pro were from accepting hallucinations. For questions on position, further conditioned on correctly identifying conditions/findings, Gemini Pro had a lower accuracy of 26.4%, with 76.68% of its errors due to accepting hallucinated positions.
Table 6: Error Analysis of GPT-4V and Gemini Pro on ProbMed. The table shows the accuracy and types of errors for three specialized question types: Abnormality, Condition/Finding, and Position. Errors are categorized into wrong answers, rejection to answer, denying ground truth, and accepting hallucinations, providing a detailed breakdown of model performance and failure modes.
Question Type Accuracy and Error Type Models
GPT-4V Gemini Pro
Abnormality Accuracy 66.06 67.05
E_wrong_answer 67.47 100.00
E_reject_to_answer 32.52 0.00
Condition/Finding Accuracy 36.90 39.97
E_deny_ground-truth 51.69 39.04
E_accept_hallucination 42.12 59.69
E_reject_to_answer 6.18 1.26
Position Accuracy 39.97 26.40
E_deny_ground-truth 39.04 23.31
E_accept_hallucination 59.69 76.68
E_reject_to_answer 1.26 0.00
4.2.3 Transferability of Domain Expertise
We conducted a finer-grained analysis to explore whether the model’s expertise in identifying features of a particular organ can be transferred to other imaging modalities. As shown in Table 16, CheXagent, a model trained exclusively on chest X-rays images, performs best in detecting abnormalities and identifying conditions/findings among all seven models when tested on chest X-ray images. We analyzed its performance to explore the transferability of expertise across the rest modalities.
As illustrated in Figure 4, CheXagent achieves significantly higher accuracy in identifying chest-related features compared to other organs, confirming our assumption that the model’s pre-training on chest X-rays enhances its performance on chest images across different modalities. Interestingly, CheXagent also demonstrated higher accuracy in identifying conditions and findings in CT scans and MRIs of the chest, achieving a 3% increase in accuracy for MRIs and a 4% increase for CT scans compared with other organs within the same unseen modality. This indicates that specialized knowledge gained on chest X-rays can be transferred to other imaging modalities of the same organ in a zero-shot manner, highlighting the potential for cross-modality expertise transfer in real-life medical imaging diagnostics.
Figure 4: Accuracy comparison of CheXagent in identifying organs and conditions/findings across different modalities. The model demonstrates significantly higher accuracy in identifying organs on chest images compared to images of other organs for both MRI and CT scans. Additionally, CheXagent shows improved accuracy in identifying conditions/findings on chest images, indicating the transferability of its specialized knowledge from chest X-ray training to other imaging modalities.
5 Conclusion
Evaluating the reliability of LMMs in the medical domain requires robust methods, and ProbMed, our newly introduced dataset, addresses this by incorporating probing evaluation and procedural diagnosis. Our study reveals significant limitations in models like GPT-4V and Gemini Pro, which perform worse than random guessing on specialized diagnostic questions, while CheXagent’s results highlight the importance of domain-specific knowledge. These findings underscore the urgent need for robust evaluation methodologies to ensure the reliability and accuracy of LMMs in medical imaging. Despite the contributions, limitations such as the imbalanced image distribution favoring Chest X-rays (see Table  2) and the lack of open-ended evaluations like report generation remain. The broader impact of our work includes improved diagnostic accuracy and patient care, but also highlights the risks of deploying unreliable models in healthcare. To mitigate these risks, we recommend rigorous testing, continuous performance monitoring, and incorporation of domain-specific expertise. Our work aims to enhance the development of trustworthy AI systems in healthcare, ultimately improving patient outcomes.
Appendix A Breakdown Results on Different Image Modality and Organ.
A.1 Brain CT Scan
Table 7: Results of different models on Brain CT scan in ProbMed. The best-performing model in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice acc 25 25 50 35.28 35.01
acc w.o. adv. pair 50 50 / 40.6 40.4
LLaVA-v1-7B acc 25.18 52.59 50 0 0
acc w.o. adv. pair 99.62 87.03 / 92.96 98.94
LLaVA-v1.6-7B acc 10.74 72.22 23.52 0 0.52
acc w.o. adv. pair 99.62 75.18 / 55.18 77.36
LLaVA-Med acc 4.81 10.74 8.82 11.85 3.15
acc w.o. adv. pair 95.55 37.4 / 26.66 92.63
MiniGPT-v2 acc 1.11 92.59 50 17.77 8.42
acc w.o. adv. pair 100 98.88 / 64.81 89.47
CheXagent acc 11.85 0 47.05 12.96 5.26
acc w.o. adv. pair 33.7 0 / 16.66 14.21
GPT-4V acc 94.07 84.07 61.76 37.03 31.05
acc w.o. adv. pair 94.81 84.44 / 55.92 49.47
Gemini Pro acc 84.44 85.18 70.58 34.81 21.05
acc w.o. adv. pair 95.92 91.48 / 45.18 37.36
num 270 270 34 270 270
A.2 Chest CT Scan
Table 8: Results of different models on Chest CT Scan in ProbMed. The best-performing model in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice acc 25 25 50 32.69 33.76
acc w.o. adv. pair 50 50 / 38.89 39.6
LLaVA-v1-7B acc 27.55 46.35 50 0.36 0.23
acc w.o. adv. pair 99.81 99.81 / 96.35 100
LLaVA-v1.6-7B acc 2.73 76.82 50 0.54 0.46
acc w.o. adv. pair 99.08 86.67 / 76.45 87.32
LLaVA-Med acc 5.47 39.78 29.41 14.41 4.37
acc w.o. adv. pair 96.71 90.51 / 35.58 95.62
MiniGPT-v2 acc 0.54 53.28 50 10.21 3.22
acc w.o. adv. pair 99.45 96.53 / 77.55 90.09
CheXagent acc 6.75 72.99 50 18.61 7.83
acc w.o. adv. pair 19.7 99.08 / 35.76 30.41
GPT-4V acc 97.07 72.94 67.64 32.9 20.78
acc w.o. adv. pair 97.62 93.96 / 47.34 39.72
Gemini Pro acc 95.62 58.21 82.35 34.48 14.28
acc w.o. adv. pair 99.27 97.26 / 60.6 61.05
num 548 548 34 548 548
A.3 Spine CT Scan
Table 9: Results of different models on Spine CT Scan in ProbMed. The best-performing model in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice acc 25 25 50 30.85 31.06
acc w.o. adv. pair 50 50 / 36.42 36.62
LLaVA-v1-7B acc 22.98 44.82 50 0 0
acc w.o. adv. pair 98.85 91.95 / 97.7 100
LLaVA-v1.6-7B acc 4.59 72.41 0 0 1.28
acc w.o. adv. pair 98.85 85.05 / 65.51 70.51
LLaVA-Med acc 2.29 11.49 50 11.49 6.41
acc w.o. adv. pair 94.25 28.73 / 26.43 79.48
MiniGPT-v2 acc 1.14 41.37 0 12.64 5.12
acc w.o. adv. pair 100 62.06 / 68.96 78.2
CheXagent acc 4.59 27.58 50 10.34 2.56
acc w.o. adv. pair 16.09 55.17 / 27.58 33.33
GPT-4V acc 81.39 69.76 0 33.73 25.97
acc w.o. adv. pair 81.39 81.39 / 39.53 42.85
Gemini Pro acc 87.2 77.9 50 22.09 25.97
acc w.o. adv. pair 94.18 93.02 / 65.11 50.64
num 86 86 2 86 86
A.4 Abdominal CT Scan
Table 10: Results of different models on Abdominal CT Scan in ProbMed. The best-performing model in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice acc 25 25 50 35.53 37.03
acc w.o. adv. pair 50 50 / 40.86 41.88
LLaVA-v1-7B acc 26.49 54.19 50 0.53 0
acc w.o. adv. pair 100 100 / 94.54 99.83
LLaVA-v1.6-7B acc 1.86 82.82 41.42 1.06 0.66
acc w.o. adv. pair 99.86 96.4 / 70.17 87.76
LLaVA-Med acc 5.05 45 30 15.44 5.28
acc w.o. adv. pair 98 93.6 / 33.28 86.11
MiniGPT-v2 acc 0 37.15 48.57 6.12 2.14
acc w.o. adv. pair 99.2 91.34 / 83.62 90.74
CheXagent acc 25.03 38.21 52.85 15.57 6.61
acc w.o. adv. pair 56.99 80.29 / 51.26 55.53
GPT-4V acc 95.72 72.72 45.71 27 23.25
acc w.o. adv. pair 96.39 92.11 / 40.9 34.71
Gemini Pro acc 98.31 69.19 65.71 28.79 20.39
acc w.o. adv. pair 99.2 96.8 / 61.86 58.2
num 750 750 70 750 750
A.5 Brain MRI
Table 11: Results of different models on Brain MRI in ProbMed. The best-performing model in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice acc 25 25 50 36.7 36.64
acc w.o. adv. pair 50 50 / 41.47 41.49
LLaVA-v1-7B acc 1.23 32.86 50 0.53 0
acc w.o. adv. pair 97.17 97.34 / 94.52 99.75
LLaVA-v1.6-7B acc 17.49 88.51 28.57 0.53 0.48
acc w.o. adv. pair 99.49 89.57 / 54.77 67.63
LLaVA-Med acc 3 8.12 23.21 14.66 2.91
acc w.o. adv. pair 90.45 44.34 / 22.08 89.29
MiniGPT-v2 acc 1.94 96.64 50 15.72 4.37
acc w.o. adv. pair 99.82 100 / 66.78 88.8
CheXagent acc 0 0 50 10.77 6.81
acc w.o. adv. pair 1.41 0 / 16.07 23.66
GPT-4V acc 96.99 94.33 58.92 36.1 27.8
acc w.o. adv. pair 97.52 95.39 / 54.69 43.17
Gemini Pro acc 95.22 94.87 78.57 35.51 19.7
acc w.o. adv. pair 96.28 99.29 / 56.89 37.71
num 566 566 56 566 566
A.6 Chest MRI
Table 12: Results of different models on Chest MRI in ProbMed. The best-performing model in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice acc 25 25 50 34.18 34.11
acc w.o. adv. pair 50 50 / 39.84 39.88
LLaVA-v1-7B acc 0 35 50 0 0
acc w.o. adv. pair 80 92.5 / 90 100
LLaVA-v1.6-7B acc 5 32.5 37.5 0 0
acc w.o. adv. pair 90 50 / 50 76.47
LLaVA-Med acc 5 45 12.5 12.5 5.88
acc w.o. adv. pair 82.5 82.5 / 17.5 82.35
MiniGPT-v2 acc 0 35 50 10 8.82
acc w.o. adv. pair 92.5 62.5 / 70 79.41
CheXagent acc 0 75 50 15 0
acc w.o. adv. pair 5 100 / 27.5 23.52
GPT-4V acc 76.92 51.28 37.5 25.64 18.18
acc w.o. adv. pair 76.92 53.84 / 30.76 30.3
Gemini Pro acc 87.5 62.5 37.5 17.5 11.76
acc w.o. adv. pair 87.5 82.5 / 55 44.11
num 40 40 8 40 40
A.7 Spine MRI
Table 13: Results of different models on Spine MRI in ProbMed. The best-performing model in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice acc 25 25 50 31.51 31.52
acc w.o. adv. pair 50 50 / 37.55 37.66
LLaVA-v1-7B acc 0 32.09 50 50 0.3
acc w.o. adv. pair 98.45 98.76 / 93.51 99.32
LLaVA-v1.6-7B acc 3.08 86.72 27.77 0.3 0
acc w.o. adv. pair 99.69 95.37 / 50.3 57.77
LLaVA-Med acc 1.54 5.24 36.11 12.96 5.4
acc w.o. adv. pair 88.27 23.45 / 17.96 77.7
MiniGPT-v2 acc 0.3 49.69 52.77 6.79 2.02
acc w.o. adv. pair 100 84.56 / 80.86 87.83
CheXagent acc 0 13.58 47.22 15.43 2.7
acc w.o. adv. pair 1.54 27.77 / 22.53 25.33
GPT-4V acc 96.28 90.71 55.55 22.6 15.59
acc w.o. adv. pair 96.28 95.66 / 32.19 26.44
Gemini Pro acc 98.13 88.81 57.14 24.53 14.91
acc w.o. adv. pair 98.13 97.82 / 50 39.32
num 332 332 35 332 332
A.8 Abdominal MRI
Table 14: Results of different models on Abdominal MRI in ProbMed. The best-performing model in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice acc 25 25 50 37.13 38.26
acc w.o. adv. pair 50 50 / 41.81 42.23
LLaVA-v1-7B acc 0 39.28 50 2.38 0
acc w.o. adv. pair 96.42 100 / 90.47 100
LLaVA-v1.6-7B acc 2.38 73.8 35.71 1.19 0
acc w.o. adv. pair 98.8 92.85 / 63.09 80.3
LLaVA-Med acc 2.38 47.61 50 14.28 9.09
acc w.o. adv. pair 79.76 95.23 / 17.85 71.21
MiniGPT-v2 acc 0 36.9 50 8.33 4.54
acc w.o. adv. pair 98.8 91.66 / 82.14 93.93
CheXagent acc 0 26.19 50 11.9 10.6
acc w.o. adv. pair 0 59.52 / 30.95 30.3
GPT-4V acc 86.9 75 50 27.38 25.75
acc w.o. adv. pair 86.9 86.9 / 50 42.42
Gemini Pro acc 89.28 72.61 85.71 28.57 25.75
acc w.o. adv. pair 91.66 94.04 / 67.85 56.06
num 84 84 14 84 84
A.9 Brain X-ray
Table 15: Results of different models on Brain X-ray in ProbMed. The best-performing model in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice acc 25 25 50 44.77 47.08
acc w.o. adv. pair 50 50 / 46.51 48.33
LLaVA-v1-7B acc 45.56 26.58 50 0 0
acc w.o. adv. pair 100 72.15 / 96.2 100
LLaVA-v1.6-7B acc 11.39 13.92 16.66 8.86 4.44
acc w.o. adv. pair 94.93 20.25 / 72.15 91.11
LLaVA-Med acc 8.86 8.86 0 20.25 4.44
acc w.o. adv. pair 100 27.84 / 24.05 88.88
MiniGPT-v2 acc 18.98 83.54 50 18.98 17.77
acc w.o. adv. pair 98.73 88.6 / 63.29 73.33
CheXagent acc 84.81 0 50 12.65 8.88
acc w.o. adv. pair 88.6 1.26 / 21.51 22.22
GPT-4V acc 82.05 8.97 33.33 43.58 22.72
acc w.o. adv. pair 83.33 8.97 / 50 36.36
Gemini Pro acc 89.87 51.89 50 31.64 31.11
acc w.o. adv. pair 93.67 62.02 / 50.63 46.66
num 79 79 6 79 79
A.10 Chest X-ray
Table 16: Results of different models on Chest X-ray in ProbMed. The best-performing model in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice acc 25 25 50 37.59 37.08
acc w.o. adv. pair 50 50 / 42.15 41.87
LLaVA-v1-7B acc 28.75 36.57 50 0.12 0.11
acc w.o. adv. pair 99.9 99.83 / 68.52 99.96
LLaVA-v1.6-7B acc 7.11 83.97 47.94 5.89 1.52
acc w.o. adv. pair 99.51 97.53 / 53.94 94.02
LLaVA-Med acc 6.25 39.77 40.24 26.28 6.14
acc w.o. adv. pair 99 90.6 / 35.67 94.34
MiniGPT-v2 acc 4.93 94.07 50.05 18.78 11.94
acc w.o. adv. pair 99.96 99.39 / 55.19 67.49
CheXagent acc 53.68 39.64 76.59 42.75 8.9
acc w.o. adv. pair 62.59 47.53 / 58.1 36.93
GPT-4V acc 91.53 67.51 53.18 39.35 21.35
acc w.o. adv. pair 91.88 94.09 / 46.18 34.38
Gemini Pro acc 98.07 76.74 61.29 25.83 15.31
acc w.o. adv. pair 99.29 99.67 / 57.83 66.42
num 3120 3120 1948 3120 3120
A.11 Spine X-ray
Table 17: Results of different models on Spine X-ray in ProbMed. The best-performing model in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice acc 25 25 50 30.95 30.99
acc w.o. adv. pair 50 50 / 37.54 37.51
LLaVA-v1-7B acc 44.55 45.04 50 0.49 0
acc w.o. adv. pair 98.51 98.51 / 93.56 98.85
LLaVA-v1.6-7B acc 4.45 82.67 33.33 1.48 0.57
acc w.o. adv. pair 99.5 95.54 / 61.38 77.58
LLaVA-Med acc 8.41 7.92 33.33 17.82 5.74
acc w.o. adv. pair 97.02 28.21 / 25.24 78.16
MiniGPT-v2 acc 2.97 52.47 58.33 16.33 4.02
acc w.o. adv. pair 99 75.24 / 61.38 69.54
CheXagent acc 82.17 20.29 62.5 16.83 0.57
acc w.o. adv. pair 90.09 40.59 / 30.69 31.6
GPT-4V acc 85.57 72.13 47.82 29.85 18.49
acc w.o. adv. pair 85.57 81.09 / 37.13 29.47
Gemini Pro acc 95.02 70.14 70.83 17.91 19.07
acc w.o. adv. pair 97.01 93.53 / 79.1 58.95
num 201 201 24 201 201
A.12 Abdominal X-ray
Table 18: Results of different models on Abdominal X-ray in ProbMed. The best-performing model in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice acc 25 25 50 36.55 37.46
acc w.o. adv. pair 50 50 / 37.54 37.51
LLaVA-v1-7B acc 53.87 53.01 50 2.15 0.56
acc w.o. adv. pair 100 99.56 / 95.25 99.45
LLaVA-v1.6-7B acc 5.17 56.46 46 6.46 1.12
acc w.o. adv. pair 97.84 78.44 / 80.17 91.52
LLaVA-Med acc 7.75 42.24 60 14.65 4.51
acc w.o. adv. pair 98.27 91.37 / 23.7 89.83
MiniGPT-v2 acc 4.74 38.79 50 18.53 5.64
acc w.o. adv. pair 99.13 86.63 / 56.03 67.23
CheXagent acc 77.15 23.7 70 12.93 2.25
acc w.o. adv. pair 90.51 46.55 / 38.79 37.85
GPT-4V acc 84.84 50.21 60 31.16 23.16
acc w.o. adv. pair 84.84 71.86 / 42.85 37.85
Gemini Pro acc 97.14 63.36 85 27.15 19.2
acc w.o. adv. pair 99.13 96.55 / 73.27 67.23
num 232 232 20 232 232
Appendix B ProbMed
Table 19: Number of questions across each question type for each image. Ground-truth questions were created based on available metadata, with "yes" answers. For each ground-truth question, we also created a corresponding adversarial question by selecting random adversarial entities and assigning "no" answers. For an image showing a normal organ without abnormality, since there is no ground-truth information on the existence of the condition and position, we only construct hallucinated questions for the condition/finding question type. For an image showing abnormality, the number of question pairs per category equals the number of existing conditions or positions.
Question type
Image with
Normal Organ
Image with
Abnormality
Modality 2 2
Organ 2 2
Abnormality 1 1
Condition/Finding 1 2 x number of existing conditions
Position 0 2 x number of existing positions